Taking a decibel to be an increase in pressure of 12.2%, by what factor has the pressure changed when the sound level has fallen by 9.00 decibels?
Answer #1

sound pressure level, SPL = 20 Log (1) SPL = 9 dB. 9 = 20 log (P/P) 9 = log ( Ple) 0.us = log (P/Po) 2.8183 - £ change in pre